calculate the percent ethanol of the concentration of a solution of acetone (CH 3 COCH 3 ) in ethanol (C 2 H 5 OH) is 2.28 molar,

Density of acetone = 0.788 g/cm 3 ; density of ethanol = 0.789 g/cm 3 .)




1
Expert's answer
2021-03-08T05:52:27-0500

2.28 moles in 1 L of Ethanol

2.28×58 gram of acetone = 132.24g

density of ethanol = 0.789

1 L ethanol = 0.789×1000 = 789 g

% ethanol = (798/921.24)×100 = 85.6 %